OpenStudy (anonymous):

Find the vertex angle of an isosceles triangle with base angles of (3x+10) and (8x-15). Show your work for credit.

OpenStudy (anonymous):

(3x+10) = (8x-15) 3x = 8x - 25 -5x = -25 x = 5